9. A diver comes off a board with arms straight up and legs straight down, giving her a moment of inertia about her rotation axis of 18 kgm2. She then tucks into a small ball, decreasing this moment of inertia to 3.6 kgm2. While tucked, she makes two complete revolutions in 1 sec. If she had not tucked at all, how many revolutions would she have made in the 1.5 sec from board to water?

10. A skater has a moment of inertia of 3 kgm2 when her arms are stretched out and 1kgm2 when her arms are brought to her sides. She starts to spin at the rate of 1 rev/sec when her arms are outstretched, and then pulls her arms to her sides. a. What is the final angular velocity? b. How much work did she have to do?
1
Expert's answer
2020-03-16T13:06:23-0400

9) We need to apply conservation of angular momentum to the diver.

The number of revolutions she makes in a certain time is proportional to her angular

velocity. The ratio of her untucked to tucked angular velocity is 


"\frac{3.6}{18}"

If she had tucked, she would have made 


"2\frac{3.6}{18}=0.4\ rev"

in the last 1.0 s, so she would have made


"0.4\frac{1.5}{1}=0.60\ rev"

  in the total 1.5 s.

10)a) From the conservation of energy:


"I_1\omega_1=I_2\omega_2"

"(3)(1)=(1)\omega_2"

"\omega_2=3\frac{rev}{s}"

b)


"W=K_2-K_1=0.5(I_2\omega_2^2-I_1\omega_1^2)"

"W=0.5((1)(3\cdot 2\pi )^2-3(2\pi )^2)=118\ J"